# Binary Classification of Light and Dark Time Traces of a Transition Edge Sensor Using Convolutional Neural Networks

### v1

@@ Abstract
The Any Light Particle Search II (ALPS II) is a light shining through a wall experiment probing the existence of axions and axion-like particles using a 1064 nm laser source. While ALPS II is already taking data using a heterodyne based detection scheme, cryogenic transition edge sensor (TES) based single-photon detectors are planned to expand the detection system for cross-checking the potential signals, for which a sensitivity on the order of $10^{-24}$ W is required. In order to reach this goal, we have investigated the use of convolutional neural networks (CNN) as binary classifiers to distinguish the experimentally measured 1064 nm photon triggered (light) pulses from background (dark) pulses. Despite extensive hyperparameter optimization, the CNN based binary classifier did not outperform our previously optimized cut-based analysis in terms of detection significance. This suggests that the used approach is not generally suitable for background suppression and improving the energy resolution of the TES. We partly attribute this to the training confusion induced by near-1064 nm black-body photon triggers in the background, which we identified as the limiting background source as concluded in our previous works. However, we argue that the problem ultimately lies in the binary classification based approach and believe that regression models would be better suitable for addressing the energy resolution. Unsupervised machine learning models, in particular neural network based autoencoders, should also be considered potential candidates for the suppression of noise in time traces. While the presented results and associated conclusions are obtained for TES designed to be used in the ALPS II experiment, they should hold equivalently well for any device whose output signal can be considered as a univariate time trace.
I Introduction
@@
Transition edge sensors (TES) are superconducting devices commonly used as microcalorimeters and bolometers when voltage-biased within the region of the superconducting phase transition [1]. In this narrow region, the resistance of the circuit changes steeply with temperature and consequently the absorption of a single photons at suitable wavelengths heat the circuit sufficiently to cause a significant change in its current. These perturbations can be efficiently detected by superconducting quantum interference devices (SQUIDs) being inductively coupled to the biased circuit. Unlike many other single-photon detectors, such as superconducting nanowire single-photon detectors (SNSPDs), TESs are capable of measuring the energy of the absorbed photons over a wide range of wavelengths. Their energy resolution together with high quantum efficiency and microsecond-scale dead time [2, 3, 4] make TESs important tools widely used in quantum computing [5, 6, 7, 8, 9], space and astrophysics experiments [10, 11, 12, 13, 14] along with particle physics and dark matter searches [15, 16, 17].
@@
A TES is planned to be used in a future science run of the Any Light Particle Search II (ALPS II) at DESY Hamburg (Germany) [18]. ALPS II aims to probe the existence of axions and axion-like particles (ALPs) [19, 20] and is essentially a light shining through a wall experiment featuring a powerful 1064 nm laser beam that is shone into a 106 m long resonant production cavity that is in 5.3 T magnetic field. While the propagation of the light is blocked by an opaque wall, the theoretically proposed photon–axion oscillation enables 1064 nm photons to emerge on the other side of the optical barrier in a regeneration cavity (symmetrical to the production cavity) [21, 22]. The detection of these reconverted photons requires an extremely sensitive detection scheme achievable with TESs [19, 20]. The target sensitivity for the conversion rate lies in the range of $10^{-5}$ Hz (one photon per day) setting the upper limit for the background rate required for the statistically significant detection of axions and ALPs within the practically feasible 20 days measurement time [19].
@@
In the recent years machine learning (ML) methods have been recognized as extremely useful tools in various fields of physics [23]. ML approaches have been widely implemented for various tasks associated with particle physics experiments, including high level data analysis, jet tagging and data simulation [24, 25, 26, 27, 28, 29, 30, 31, 32]. Lately, ML methods have also been used to detect hints from axion-like particles in LHC [33] and pulsar dispersion measurements [34], and are planned to be implemented for analyzing data in the SuperCDMS direct dark matter experiment [35]. Most relevant for our interest, Manenti et al. [36] have recently applied unsupervised ML models to study the background of a TES system not connected to the outside of a cryostat, by e.g. a fiber (intrinsic background). They report that the majority of the observed background pulses originate from high-energy events associated with radioactive decays and secondary cosmic-ray particles. The resulting dark counts are rather easy to distinguish from low-energy photon induced pulses by simply comparing the pulse shape, as already concluded in our previous work [37]. This is because the energy released from the various high-energy events is likely to be deposited within the underlying substrate rather than the TES itself. Due to slow diffusion of heat from the substrate to the TES, the dark counts have significantly larger rise and decay times when compared with typical photon induced pulses where the energy is deposited directly to the TES [36, 38].
@@
While the unsupervised ML models have been mainly used for qualitative categorization of the recorded pulses, supervised ML models are better suited for actual quantitative background suppression. One can expect the state-of-the-art supervised ML models to outperform the capabilities of traditional data processing techniques. We have successfully implemented this in the past for our recorded intrinsic background without an optical fiber connected to the TES [37]. In this work, we expand on this study by also considering the extrinsic background measured while an optical fiber links the lab environment to the TES inside the dilution refrigerator. This mainly introduces an additional background resulting from fiber coupled black-body radiation. The black-body photons have been identified as the limiting background for our experimental setup [20, 39, 40], and have been previously addressed using a traditional cut-based analysis without relying on machine learning [38]. We want to point out that the black-body background rate is ultimately determined by the energy resolution of the TES since lower energy resolution enables more reliable distinction between the signal and the background photons. Thus, different analysis methods addressing the suppression of noise in the time traces differently can have significant effects on the background rates [40]. For example, we have found that performing fits to the measured TES pulses in frequency domain instead of time domain results in 2-fold improvement in the energy resolution [39, 38].
@@
Figure 1: (a) The average measured light pulse and dark pulses, where the shaded regions represent the associated standard deviations. The inset presents a randomly chosen light pulse as an example of the signal and noise shapes. (b) Principal Component Analysis (PCA) scatter plot showing the projection of pulse feature vectors ($\tau_{\mathrm{rise}}$, $\tau_{\mathrm{decay}}$, $\chi_{\mathrm{Ph}}^{2}$, $\mathrm{Ph}_{\mathrm{FFT}}$, $\chi^{2}_{\mathrm{FFT}}$) onto the first two principal components (PC1 and PC2). The inset shows a close-up of the cluster associated with light pulses, showing overlap with some of the dark pulses measured in extrinsic background.
@@
Here, for the first time, we are trying to further improve the rejection of near-1064 nm black-body photons using convolutional neural networks (CNN) that are considered as the state-of-the-art machine learning model for univariate time-series classification. Ultimately, the goal is to reach a background rate below $10^{-5}\,\mathrm{Hz}$ while maintaining a tolerable rate for correctly classified signal pulses (analysis efficiency). The CNNs expand the architecture of the conventional multi-layer perceptrons (feedforward neural networks) via the introduction of convolutional layers that apply different filters (kernels) to the input data enabling the efficient extraction of spatial (or temporal) patterns [41]. The CNNs remain the underlying technology behind the state-of-the-art image classification ML-models [42, 43, 41], also covering the univariate time series classifiers [44, 45, 46, 47]. Consequently, CNNs are expected to show the best performance in the suppression of the background triggers. A major benefit of CNNs is that they enable model-independent analysis of recorded pulses as one does not have to rely on fitting any specific functions to the data. We will utilize the CNNs as binary classifiers that are trained to distinguish between 1064 nm photon induced light pulses and any other background source induced dark pulses. These classifiers are then ensembled to quantitatively study the background sources that are particularly difficult to distinguish from the light pulses and to see whether the CNNs can be used for further background suppression.
@@
The manuscript is organized as follows: in Section II we describe our experimental setup and how the experimental data used in this work was obtained (see Ref. [38] for more details). Next, in Section III we present a detailed description of the overall architecture of the considered CNN and explain how we use the experimentally measured data to train it and evaluate its performance. Details of the CNN’s hyperparameter optimization are also presented. The performance of the optimized CNN is analyzed and further fine-tuned in Section IV. We then proceed in evaluating the true performance of the model in classifying the light and dark pulses in Section V and study the observed background pulses in detail in Section VI. Finally, we discuss the background source induced confusion in the CNN training process in Section VII before summarizing the final conclusions in Section VIII.
II Experimental data
@@
All of the experimental data was measured using a tungsten-based TES chip fabricated at National Institute of Standards and Technology (NIST). The TES has been optimized for the detection of 1064 nm photons through its multilayer structure [48]. The working point of the voltage-biased TES was set to 30% of its normal state resistance and its current was monitored by an inductively coupled SQUID (manufactured by Physikalisch Technische Bundesanstalt; PTB) with 5 GHz gain bandwidth product at a sampling rate of 50 MHz. The photons absorbed by the TES were then detected as pulses in the SQUID output voltage $V_{\mathrm{out}}(t)$. The TES+SQUID module was operated within a Bluefors SD dilution refrigerator at 25 mK base temperature.
@@
describing a photonic event triggered TES pulse at around $t=t_{0}-\tau_{\mathrm{rise}}$. The shape of the pulse is determined by the rise and decay times $\tau_{\mathrm{rise}}$ and $\tau_{\mathrm{decay}}$, respectively, the obtained distribution of which will be used to determine the cuts. The $\chi_{\mathrm{Ph}}^{2}$-error associated with the performed fit is also considered. In addition, we have also considered a fitting function from Small Signal Theory [1]
@@
While fitting Eq. (2) in time domain is unstable, we have performed the fit in frequency domain, in particular due to the simple Fourier transformation of Eq. (2). The obtained fitting parameters were then used to calculate the associated peak height ($\mathrm{Ph_{\mathrm{FFT}}}$) which will also be considered for the filtering of the pulses. This specific parameter was chosen because its associated distribution for light pulses has previously resulted in the lowest achieved energy resolution for our TES [39, 38]. The $\chi_{\mathrm{FFT}}^{2}$-error associated with the fits in frequency domain is also considered for filtering.
In order to mitigate the effects of scattering processes and nonlinear optical effects (that can alter the wavelength of the photons emitted from the laser) on the training process, we have subjected the 3928 single-photon triggers for minor filtering. This was done by only including pulses whose $\tau_{\mathrm{rise}}$, $\tau_{\mathrm{decay}}$ and $\mathrm{Ph}_{\mathrm{FFT}}$ were simultaneously within 0.1%–99.9% quantiles of their associated distributions while $\chi_{\mathrm{Ph}}^{2}$ and $\chi^{2}_{\mathrm{FFT}}$ being below 99.9% quantiles. This resulted in the discrimination of 0.76% (30) of the filtered triggers, leaving us a total of 3898 pulses that are used for training and evaluating the CNNs.
@@
Figure 2: A schematic illustration of the basic architecture of the considered CNN and its hyperparameters whose optimization is explicitly addressed (see Table I).
In summary, after data cleaning we are left with 3898 light pulses and 8872 dark pulses, making the overall size of the dataset 12,770 to be used for the training and evaluation the CNNs. Before proceeding, we want to further characterize the differences between light and dark pulses via Principal Components Analysis (PCA) in order to detect any possible overlap between the resulting light and dark clusters indicating the presence of photonic background. We have done this by associating each pulse with a feature vector assembled from the above introduced fitting parameters as ($\tau_{\mathrm{rise}}$, $\tau_{\mathrm{decay}}$, $\chi_{\mathrm{Ph}}^{2}$, $\mathrm{Ph}_{\mathrm{FFT}}$, $\chi^{2}_{\mathrm{FFT}}$). The PCA scatter plot visualizing the projection of these feature vectors for both light and dark pulses onto the two main principal components is presented in Fig. 1(b). As expected, the light pulses are tightly clustered in one spot while the dark pulses are much more spread out. Regardless, one can observe significant overlap between the light and dark pulses as illustrated in the inset of Fig. 1(b), most likely originating from fiber coupled black-body radiation [40]. We will analyze this later in the paper aided by the CNN ensembles.
III CNN Architecture
The basic architecture of the CNN considered in this manuscript resembles the generally used structure for image classifying tasks [49, 42]. As illustrated in Fig. 2, the CNN consists of pairs of i) convolutional and average pooling layers followed by ii) flatten and dropout layers connected to iii) dense layers with gradually decreasing sizes ultimately ending up to a single output neuron. As typical for binary classifiers, we use binary cross-entropy (log-loss) as the loss function to guide the training process. The weights of the model are initialized using the Glorot uniform approach [50] and updated during training following the Adaptive Moment Estimation (Adam) algorithm [51].
In order to limit the size of the search space for hyperparameter optimization, we fix the activation functions associated with the convolutional layers to tanh while using ReLu for the dense layers. This combination was observed to clearly result in best performance in our initial tests of the model. We further require that all of the convolutional layers share the same number of filters, filter size, and pool size. We fix the pool size to 2, limiting the maximum number of convolutional layers in the considered architecture to 10. The structure of the dense layers is limited by requiring that the number of neurons must always drop by half in the following hidden layer. This leaves only the maximum number of neurons within the first layer and the number of hidden layers as the only hyperparameters to be optimized for the dense part of the CNN. On top of the architectural hyperparameters, we address the optimization of the dropout rate, number of epochs and batch size. A summary of the considered search space for the hyperparameter optimization is presented in Table I. The CNN is implemented using a high-level neural network API Keras version 2.12.0 [52] with a TensorFlow version 2.12.1 backend [53].

### v2

# Binary Classification of Signal and Background Triggers of a Transition Edge Sensor Using Convolutional Neural Networks

@@ Abstract
The Any Light Particle Search II (ALPS II) is a light shining through a wall experiment probing the existence of axions and axion-like particles using a 1064 nm laser source. While ALPS II is already taking data using a heterodyne based detection scheme, cryogenic transition edge sensor (TES) based single-photon detectors are planned to expand the detection system for cross-checking the potential signals, for which a sensitivity on the order of 10-24 W is required. In order to reach this goal, we have investigated the use of convolutional neural networks (CNN) as binary classifiers to distinguish the experimentally measured 1064 nm photon triggered (light) pulses from background (dark) pulses. Despite rigorous hyperparameter optimization, the CNN based binary classifier did not outperform our previously optimized cut-based analysis in terms of detection significance. Our findings suggest that training confusion, introduced by near-1064 nm black-body photon triggers in the extrinsics background, is a significant factor limiting the CNNs performance for the associated dataset. The fiber coupled black-body radiation was identified as the limiting background source as concluded in our previous works. Given our results, we recommend that future studies explore regression-based CNNs, placing greater emphasis on the use of standardized and carefully structured training data rather than on extensive hyperparameter optimization. While the presented results and associated conclusions are obtained for TES designed to be used in the ALPS II experiment, they should hold equivalently well for any device whose output signal can be considered as a univariate time trace.
1 Introduction
@@
Transition edge sensors (TES) are superconducting microcalorimeters that are voltage-biased within the region of superconducting phase transition where the resistance of the TES changes steeply with temperature [Irwin2005transition]. Here, the absorption of a single photon heats up the TES sufficiently to result in a significant change in its bias current. These current perturbations are detected by an iductively coupled superconducting quantum interference device (SQUID). Unlike many other single-photon detectors, such as superconducting nanowire single-photon detectors (SNSPDs), TESs are capable of measuring the energy of the absorbed photons over a wide range of wavelengths. Their energy resolution together with high quantum efficiency and microsecond-scale dead time [Lita2008counting, Hattori2022optical, DeLucia2024transition] make TESs important tools widely used in quantum computing [Lita2010superconducting, Karasik2011nanobolometers, Mattioli2015photon, Hummatov2023fast, Li2023multi], space and astrophysics experiments [Romani1999first, Bruijn2003development, Goldie2011ultra, Bergen2016design, Appel2022calibration] along with particle physics and dark matter searches [Miyazaki2020dark, Angloher2024doubletes, Romani2024transition].
@@
A TES is planned to be used in a future science run of the Any Light Particle Search II (ALPS II) at DESY Hamburg (Germany) [Bahre2013any]. ALPS II aims to probe the existence of axions and axion-like particles (ALPs) [Shah2021tes, Gimeno2023tes] and is essentially a light shining through a wall experiment featuring a powerful 1064 nm laser beam that is shone into a 106 m long resonant production cavity that is in 5.3 T magnetic field. While the propagation of the light is blocked by an opaque wall, the theoretically proposed photon–axion oscillation enables 1064 nm photons to emerge on the other side of the optical barrier in a regeneration cavity (symmetrical to the production cavity) [Sikivie1983experimental, Isleif2022any]. The detection of these reconverted photons requires an extremely sensitive detection scheme achievable with TESs [Shah2021tes, Gimeno2023tes]. The target sensitivity for the conversion rate lies in the range of $10^{-5}$ Hz (one photon per day) setting the upper limit for the background rate required for the statistically significant detection of axions and ALPs within the practically feasible 20 days measurement time [Shah2021tes].
@@
In the recent years machine learning (ML) methods have been recognized as useful tools in various fields of physics [Carleo2019machine]. ML approaches have been widely implemented for various tasks associated with particle physics experiments, including high level data analysis, jet tagging and data simulation [Dery2017weakly, Barnard2017parton, Collado2021learning, Du2021deep, Graczykowski2022using]. ML based analysis of time traces, in particular, has been widely implemented in nuclear spectroscopy [Zehtabvar2024review], where neural networks have demonstrated significant performance in noise suppression [Regadio2019filtering]. Lately, ML methods have also been used to detect hints from axion-like particles in LHC [Ren2021detecting] and pulsar dispersion measurements [Shi2025machine], and are planned to be implemented for analyzing data in the SuperCDMS direct dark matter experiment [Cushman2024strategies]. Most relevant for our interest, Manenti et al. [Manenti2024dark] have recently applied unsupervised ML models to study the background of a TES system that is not connected to the outside of a cryostat by an optical fiber (intrinsic background). They report that the majority of the observed background pulses originate from high-energy events associated with radioactive decays and secondary cosmic-ray particles. The resulting dark counts are easy to distinguish from low-energy photon triggers by simply comparing the pulse shapes, as already concluded in our previous work [Meyer2024first]. This is because the energy released from the various high-energy events is likely to be deposited within the underlying substrate rather than the TES itself. Due to slow diffusion of heat from the substrate to the TES, the intrinsic dark counts have significantly larger rise and decay times when compared with typical photon induced pulses where the energy is deposited directly to the TES [Manenti2024dark, RubieraGimeno2024optimizing].
@@
While the unsupervised ML models have been mainly used for qualitative categorization of the recorded pulses, supervised ML models are better suited for actual quantitative background suppression. One can expect the state-of-the-art supervised ML models to outperform the capabilities of traditional data processing techniques. We have successfully implemented this in the past for intrinsic background [Meyer2024first]. In this work, we expand on this study by also considering the extrinsic background measured while an optical fiber links the lab environment to the TES inside the dilution refrigerator. This mainly introduces an additional background resulting from fiber coupled black-body radiation. The black-body photons have been identified as the limiting background for our experimental setup [Gimeno2023tes, Gimeno2024tes, Gimeno2025simulation], and have been previously addressed using a traditional cut-based analysis without relying on machine learning [RubieraGimeno2024optimizing]. We want to point out that the black-body background rate is ultimately determined by the energy resolution of the TES since higher energy resolution (smaller quantitative value) enables more reliable distinction between the signal and the background photons. Thus, different analysis methods addressing noise suppression differently can have significant effects on the background rates [Gimeno2025simulation]. For example, we have previously found that performing fits to the measured TES pulses in frequency domain instead of time domain results in 2-fold improvement in the energy resolution [Gimeno2024tes, RubieraGimeno2024optimizing].
Figure 1: (a) A circuit diagram of the experimental TES setup. Applying the constant bias voltage ($V_{\mathrm{bias}}$) and the shunt resistor ($R_{\mathrm{shunt}}$) parallel to the TES allows to set the TES working point (negative electrothermal feedback). (b) A schematic illustration of the layer structure of the TES optimized for detecting 1064 nm photons, where the 20 nm thick superconducting W layer with a surface area of $25\,\mathrm{\mu m}\times 25\,\mathrm{\mu m}$ acts as the active material. (c) Picture of the used TES+SQUID module, including two TESs and their associated SQUIDs.
@@
Figure 2: (a) The average measured light pulse and dark pulses, where the shaded regions represent the associated standard deviations. The inset presents a randomly chosen light pulse as an example of the signal and noise shapes. (b) Principal Component Analysis (PCA) scatter plot showing the projection of pulse feature vectors ($\tau_{\mathrm{rise}}$, $\tau_{\mathrm{decay}}$, $\chi_{\mathrm{ph}}^{2}$, $\mathrm{V}_{\mathrm{min,\,FFT}}$, $\chi^{2}_{\mathrm{FFT}}$) onto the first two principal components (PC1 and PC2). The inset shows a close-up of the cluster associated with light pulses, showing overlap with some of the dark pulses measured in extrinsic background.
@@
Here, for the first time, we are trying to further improve the rejection of near-1064 nm black-body photons using convolutional neural networks (CNN) that are considered as the state-of-the-art machine learning model for univariate time-series classification. Ultimately, the goal is to reach a background rate below $10^{-5}\,\mathrm{Hz}$ while maintaining a tolerable rate for correctly classified signal pulses (analysis efficiency). The CNNs expand the architecture of the conventional multi-layer perceptrons (feedforward neural networks) via the introduction of convolutional layers that apply different filters (kernels) to the input data enabling the efficient extraction of spatial (or temporal) patterns [Li2021survey]. The CNNs remain the underlying technology behind the state-of-the-art image classification ML models [Lecun2015deep, Rawat2017deep, Li2021survey], also covering the univariate time series classifiers [Foumani2023deep, Wang2017time, Ismail2020inceptiontime, Dempster2020rocket]. Consequently, CNNs are expected to show the best performance in the suppression of the background triggers. A major benefit of CNNs is that they enable model-independent analysis of recorded pulses as one does not have to rely on fitting functions to the data. We will utilize the CNNs as binary classifiers that are trained to distinguish between 1064 nm photon induced light pulses and any other background source induced dark pulses. These classifiers are then ensembled to quantitatively study the background sources that are particularly difficult to distinguish from the light pulses and to see whether the CNNs can be used for further background suppression.
@@
The manuscript is organized as follows: in Section 2 we describe our experimental setup and how the experimental data used in this work was obtained (see Ref. [RubieraGimeno2024optimizing] for more details). Next, in Section 3 we present a detailed description of the overall architecture of the considered CNN and explain how we use the experimentally measured data to train it and evaluate its performance. Details of the CNN’s hyperparameter optimization are also presented. The performance of the optimized CNN is analyzed and further fine-tuned in Section 4. We then proceed in evaluating the true performance of the model in classifying the light and dark pulses in Section 5 and study the observed background pulses in detail in Section 6. Finally, we discuss the background source induced confusion in the CNN training process in Section 7 before summarizing the final conclusions in Section 8.
2 Experimental data
@@
All of the experimental data was measured using a tungsten-based TES chip fabricated at National Institute of Standards and Technology (NIST). The working point of the voltage-biased TES was set to 30% of its normal state resistance and its current was monitored by an inductively coupled flux-locked SQUID (manufactured by Physikalisch Technische Bundesanstalt; PTB) with 5 GHz gain bandwidth product at a sampling rate of 50 MHz. The photons absorbed by the TES were then detected as pulses in the SQUID output voltage $V_{\mathrm{out}}(t)$. The TES+SQUID module was operated within a Bluefors SD dilution refrigerator at 25 mK base temperature. The effective circuit diagram of the used TES setup is illustrated in Fig. 1(a), while the layer structure of the TES, optimized for the detection of 1064 nm photons [Shah2022characterising], is presented in Fig. 1(b). A picture of the TES+SQUID module installed into the dilution refrigerator is shown in Fig. 1(c).
@@
describing a photonic event triggered TES pulse at around $t=t_{0}-\tau_{\mathrm{rise}}$. The parameter $A_{\mathrm{ph}}$ is directly proportional to the amplitude of the pulse, while the pulse shape is determined by the rise and decay times $\tau_{\mathrm{rise}}$ and $\tau_{\mathrm{decay}}$, respectively. The obtained distribution of $\tau_{\mathrm{rise}}$ and $\tau_{\mathrm{decay}}$ will be used in determining the cuts. The $\chi_{\mathrm{ph}}^{2}$-error associated with the performed fit is also considered. In addition, we have also considered a fitting function from Small Signal Theory [Irwin2005transition]
@@
where the parameters $A_{\mathrm{FFT}}$, $\tau_{+}$ and $\tau_{-}$ are analogous to the $A_{\mathrm{ph}}$, $\tau_{\mathrm{rise}}$ and $\tau_{\mathrm{decay}}$ of the phenomenological model introduced above. While fitting Eq. (2) in time domain is unstable, we have performed the fit in frequency domain, in particular due to the simple Fourier transformation of Eq. (2). The obtained fitting parameters were then used to calculate the associated peak height ($\mathrm{V}_{\mathrm{min,\,FFT}}$) which will also be considered for the filtering of the pulses. This specific parameter was chosen because its associated distribution for light pulses has previously resulted in the highest achieved energy resolution for our TES [Gimeno2024tes, RubieraGimeno2024optimizing]. The $\chi_{\mathrm{FFT}}^{2}$-error associated with the fits in frequency domain is also considered for filtering.
In order to mitigate the effects of scattering processes and nonlinear optical effects (that can alter the wavelength of the photons emitted from the laser) on the training process, we have subjected the 3928 single-photon triggers for minor filtering. This was done by only including pulses whose $\tau_{\mathrm{rise}}$, $\tau_{\mathrm{decay}}$ and $\mathrm{V}_{\mathrm{min,\,FFT}}$ were simultaneously within 0.1%–99.9% quantiles of their associated distributions while $\chi_{\mathrm{ph}}^{2}$ and $\chi^{2}_{\mathrm{FFT}}$ being below 99.9% quantiles. This resulted in the discrimination of 0.76% (30) of the filtered triggers, leaving us a total of 3898 pulses that are used for training and evaluating the CNNs.
The filtered 3898 pulses were further post-processed by removing a possible voltage offset by shifting the recorded $V_{\mathrm{out}}(t)$ values by the average voltage value measured between $t=0\texttt{-}24\,\mu\mathrm{s}$. The waveforms were truncated to a time window of $24\,\mu\mathrm{s}$ (corresponding to 1200 samples) by locating the pulse minimum and including the 300 antecedent and 900 subsequent samples. This ensures that we fully capture the 1064 nm photon triggered pulses given their average rise and decay times (see Fig. 2(a)). Evidently, we strived for minimizing the time window (and the number of samples) in order to limit the computational load for the CNNs. The waveforms were not resampled prior to training of the CNNs in order to prevent any loss of information. For the rest of the manuscript, we will keep referring to these 1064 nm photon triggered time traces as light pulses. The average measured light pulse is presented in Fig. 2(a) together with an example of a single pulse in the inset of the figure, further illustrating the baseline noise present in the measurements.
Figure 3: A schematic illustration of the basic architecture of the considered CNN and its hyperparameters whose optimization is explicitly addressed (see Table 1).
@@
3 CNN Architecture
The basic architecture of the CNN considered in this manuscript resembles the generally used structure for image classifying tasks [Krizhevsky2012imagenet, Lecun2015deep]. As illustrated in Fig. 3, the CNN consists of pairs of i) convolutional and average pooling layers followed by ii) flatten and dropout layers connected to iii) dense layers with gradually decreasing sizes ultimately ending up to a single output neuron. As typical for binary classifiers, we use binary cross-entropy (log-loss) as the loss function to guide the training process. The weights of the model are initialized using the Glorot uniform approach [Glorot2010understanding] and updated during training following the Adaptive Moment Estimation (Adam) algorithm [Kingma2014adam].
In order to limit the size of the search space for hyperparameter optimization, we fix the activation functions associated with the convolutional layers to tanh while using ReLu for the dense layers. This combination was observed to clearly result in best performance in our initial tests of the model. We further require that all of the convolutional layers share the same number of filters, filter size, and pool size. We fix the pool size to 2, limiting the maximum number of convolutional layers in the considered architecture to 10. The structure of the dense layers is limited by requiring that the number of neurons must always drop by half in the following hidden layer. This leaves only the maximum number of neurons within the first layer and the number of hidden layers as the only hyperparameters to be optimized for the dense part of the CNN. On top of the architectural hyperparameters, we address the optimization of the dropout rate, number of epochs and batch size. A summary of the considered search space for the hyperparameter optimization is presented in Table 1. The CNN is implemented using a high-level neural network API Keras version 2.12.0 [Chollet2015keras] with a TensorFlow version 2.12.1 backend [Tensorflow2015-whitepaper].